If the federal government was required to balance its budget, when would government spending most likely decrease?

Select one:

a. When employment was rising rapidly for an extended period of time

b. When unemployment was declining for an extended period of time

c. When the economy was experiencing a strong expansion due to high business confidence

d. When the economy was experiencing a demand pulled inflation

e. A balance budget amendments ensures that government spending never decrease

Suppose that the Price level = 140, Real GDP = $15 billion and Velocity of Money = 21. What does the supply of money equal?

Select one:

a. $200 billion

b. $180 billion

c. $150 billion

d. $120 billion

e. $100 billion
Answer #1

1)

Answer: (A)

High employments means now the government needs to spend less to balance the budget. Thus, the decrease in spending would create a balanced budget.

2)

MV = PY

21*M =15*140

M =2100/21

=100

Answer: (E)